# REDS: Resource-Efficient Deep Subnetworks for Dynamic Resource Constraints

### v2

@@ 1 Introduction
Data processing pipelines in edge devices increasingly rely on deep learning models to identify patterns and extract insights from multimodal IoT data. Examples include predictive maintenance in industrial automation, object identification and tracking in smart camera systems (Qu et al., 2022), activity and healthcare trackers in mobile (Ravi et al., 2016), wearable and hearable applications (Maag et al., 2017; Sabry et al., 2022). In all these systems, deep models are deployed and run along with other tasks, under constraints and priorities dictated by the current context and available resources, including storage, CPU time, energy and bandwidth. A large body of work explore different techniques to optimize and compress deep models without hurting accuracy and generalization abilities, while accelerating their execution in software (Boehm et al., 2018) and in hardware (Jouppi et al., 2017). Model pruning and quantization (Han et al., 2016) have become part of standard deep learning deployment pipelines, e.g., TFLMicro (David et al., 2021) and TensorRT (Vanholder, 2016), to enable deep learning on severely constrained embedded hardware operated by low-power microcontrollers with only a few kB of RAM. Pruning covers a set of methods that reduce the model size by eliminating unimportant operations (weights, neurons, kernels) in the model (Dai et al., 2018; Li et al., 2017). These methods date back to the optimal brain damage (LeCun and et al., 1990) and the optimal brain surgeon (Hassibi and Stork, 1993), which suggest to prune the weights based on the Hessians of the loss function. Recent pruning methods (Han et al., 2015; Entezari and Saukh, 2019; Timpl et al., 2022) propose to prune network weights with a low magnitude or a low magnitude increase. (Li et al., 2017) propose to prune channels in CNNs based on a filter weight norm, while Hu et al. (2016) use the average percentage of zeros in the output to prune unimportant channels.
@@
One drawback of compile-time optimizations, e.g., pruning, is that the resulting models are resource-agnostic. They thus yield suboptimal performance in many interesting applications where resource availability depends on different dynamically changing factors such as available energy, task priority and timing constraints. Another drawback of one-shot model compression techniques, is that these are applied to the whole model, making exploration of different options for a resource-aware on-device model reconfiguration challenging. Both problems are described in detail below.
@@
Dynamic resource constraints. Many interesting applications can make use of resource-aware deep models, i.e., models that can adapt their execution to available computational resources and time constraints. For example, camera image processing by a drone or a car may depend on the respective speed (Qu et al., 2022). Processing interesting and relevant data can justify using more energy and computational time than when running regular environment scans (Gherman et al., 2021). A naive solution to address dynamic resource constraints is to store several independent deep models and switch between them as resource availability and task priorities change. The drawback of this approach is both the increased memory consumption to store these independent models which does not scale, and the overhead of switching between models at runtime, e.g., by loading these from flash to RAM and reallocating the necessary tensors.
@@
Several approaches have been proposed to adapt a model to dynamic resource constraints. These can be classified into the methods that implement early exit predictions at a cost of a reduced accuracy, e.g., BudgetRNN (Kannan and Hoffmann, 2021) and ASVN (Bambusi et al., 2021), and those that build a subnetwork structure using weight sharing, with each subnetwork being more efficient yet possibly less accurate, e.g., DRESS (Qu et al., 2022) and NestDNN (Fang et al., 2018). This work falls into the latter category, yet makes use of structured sparsity constructively to allow for additional flexibility and hardware support on typical IoT devices.
@@
Deep learning support on IoT devices. Modern IoT hardware often provides support for running deep learning models. This includes support for floating point instructions, dedicated instructions for frequent operations, such as MLA, FMA, and their vector versions. Deep model frameworks for a specific platform make use of the available hardware-specific features to provide maximum speedup. However, the methods that introduce subnetwork structures to a model can not rely on the toolchain support to optimize the execution of each subnetwork. For examples, fine-grained weight pruning may lead to accelerated execution due to an abundance of zeros in the weight matrices and sparse filters, yet unconstrained locations of these zeros present a difficulty to make use of unstructured sparsity. Moreover, there are specific algorithms designed to best handle different sparsity levels (Hoefler et al., 2021), that may or may not enjoy the available hardware support. To overcome the issues, NestDNN (Fang et al., 2018) prunes convolutional filters; these have to be paged in or out when switching from one multi-capacity model to another. DRESS (Qu et al., 2022) relies on a specialized hardware to ensure the applied sparsity pattern translates into computation efficiency. Hardware accelerators may, however, appear too power-hungry and expensive for battery-powered IoT devices. Specialized hardware is also inflexible as the deep learning technology advances fast and hardware modernizations are costly.
Contributions. We present the design of Resource-Efficient Deep Subnetworks (REDS), featuring a nested submodel structure to address dynamic resource constraints on mobile and IoT devices. In contrast to DRESS and NestDNN, REDS introduces a novel hardware-aware method to design the nested subnetwork structure, i.e., the structured sparsity patterns, by formulating and solving an iterative knapsack problem, provide theoretical guarantees and empirical evidence of outstanding solution performance. Furthemore, we leverage permutation invariance of neurons (Entezari et al., 2021) to keep the subnetwork weight tensors in contiguous memory regions, i.e., dense layers remain dense in all lower-capacity subnetworks. This makes the description of the subnetwork structure elegant, reduces REDS adaptation time, and allows for further hardware-specific optimizations. REDS code will be made publicly available. [^1] Link: https://github.com/FraCorti/Deep_Subnetworks_for_Dynamic_Resource_Constraints Our contributions are:
@@
  • We formulate the optimization problem as an iterative knapsack, present theoretical analysis and provide empirical evidence of the solution effectiveness, especially in the low-data regime (Sec. 3 and Sec. 4).
  • REDS make use of the permutation invariance of neurons to enable hardware-specific optimizations (Sec. 3). In particular, for resource-constrained devices that use data caches, REDS provide a compile-time optimization to ensure all subnetwork weights reside in contiguous memory (Sec. 5).
In the next section we explain REDS on a simple fully-connected neural network (DNN).
2 A Base Example
@@
We first illustrate advantages of REDS on a fully-connected network depicted in Fig. 1. The three dense weight matrices $W_{1}^{3\times 4}$, $W_{2}^{4\times 4}$ and $W_{3}^{4\times 2}$ connect neurons in consecutive layers that are stored in memory as unfolded one-dimensional arrays. We assume the row-major format is used to store and access model weights in memory, which is the standard choice on most hardware architectures. Given a trained and optimized model, weight matrices are mapped to continuous memory regions, allowing for a straightforward use of vector instructions to speed up on-device inference.
@@
REDS organizes a model into a set of nested submodels, i.e., the active weights of a child subnetwork are fully contained in its parent subnetwork. To enable this structure, we slice each parent subnetwork into an active and an inactive part, where the active part shapes the child subnetwork. REDS makes use of structured sparsity, i.e., model slicing occurs at the level of individual neurons and convolutional filters. Slicing off two neurons in the first hidden layer in Fig. 1 leads to the removal of two columns and two rows in the weight matrices $W_{1}$ and $W_{2}$ respectively.
@@
Firstly, even though any neuron can be removed from a layer, we re-order neurons to have a group of active neurons followed by inactive neurons due to the permutation invariance phenomenon of neural networks (Entezari et al., 2021; Jordan et al., 2022). A permutation does not change the function of the network, but allows optimizing the memory layout to keep subnetwork weights in contiguous memory. The above observations make the subnetwork layout to be stored efficiently on-device. In fact, it is sufficient to store one integer value denoting the slicing point for each layer in each subnetwork, which corresponds to the number of active computational units. This is possible since active and inactive units build continuous groups in memory. Activating a particular subnetwork means changing the size of the dimension of the weight tensor in each layer. An additional bit is used to indicate optimization for caches, i.e., that a flipped operation order is applied.
@@
Secondly, REDS achieves minimal adaptation overhead: only the width of the layers has to be updated to switch to a different model. Switching from one REDS submodel to another requires adjusting only the sizes of the layers. Due to a local scope of a slice, the modification affects only the incoming and outgoing connections; inactive computational units do not participate in inference. Examples of slicing a dense and a convolutional networks are shown in Fig. 1 and Fig. 2. The computational cost of running inference using a subnetwork is not affected by the presence of other REDS networks, in sharp contrast to the approaches that use binary masks to select active neurons or channels (Qu et al., 2022; Mishra et al., 2021).
@@
Finally, pruning a neuron in one layer may remove more multiply–accumulate operations (MACs) than in another layer. In our example in Fig. 1, removing one neuron in the first hidden layer yields more MAC reduction than removing a neuron in the second hidden layer due to a different number of connections. Moreover, the contribution of these neurons to model accuracy may be different. Related research suggests several importance measures for individual weights, neurons, and convolutional filters (Shen et al., 2022; Molchanov et al., 2019; Hoefler et al., 2021; Frantar and Alistarh, 2022). These can be used to optimize a network for performance while keeping essential elements. Our problem is different: we build a nested structure, and this requires iterating over the subnetworks. DRESS starts from the largest network. NestDNN implements a top-down pruning and a bottom-up freeze-and-regrow algorithm.
We formulate the optimization problem for a parent-child subnetwork pair as a variant of the knapsack problem (Shen et al., 2022). To extend the solution to multiple nested subnetworks, we generalize the approach to an iterative knapsack problem (Della Croce et al., 2019). In contrast to Shen et al. (2022), our knapsack formulation models the dependency between layers by iteratively chaining constraints to enforce each subnetwork’s weight tensors to be functionally correct (see Fig. 2). Furthermore, REDS knapsack formulation can constraint each subnetwork’s peak memory usage known to be the major memory bottleneck for enabling neural networks on edge computing devices (Lin and et al., 2023). Our theoretical findings (Appendix A.2) suggest that the knapsack method applied to the smallest subnetwork first and letting it grow, i.e., the bottom-up approach, is more effective than starting from the largest subnetwork and pruning it down to the smallest one. We empirically show this in Sec. 4.
@@
In the following sections we focus on specific contributions and design choices for REDS on the way to a fully-functioning framework. Sec. 3 discusses a pipeline to convert a model to REDS and presents a hardware-aware iterative knapsack method to choose the model slicing points. Sec. 5 revisits the memory layout optimization for cache and provides an empirical evaluation of the achieved gain on embedded hardware.
@@ 3 Resource-Efficient Deep Subnets
Before describing the pipeline how we build REDS, we first argue why layer slicing by skipping a contiguous sequence of computational units, such as neurons or convolutional filters, doesn’t constrain the expressiveness of REDS in any way. These computational units in all but the last layer of the network can be pruned arbitrary, but can be reordered to form a contiguous sequence without changing the network function. We then describe REDS construction pipeline. In the heart of it is the novel method to decide on the choice of the subnetwork structure of REDS based on the knapsack problem. Finally, we discuss how to fine-tune the submodels in REDS to recover lost accuracy due to the applied pruning of computational units.
@@ 3.1 Permutation invariance
The optimization landscape of neural networks contains an abundance of minima and saddle points as a result of numerous symmetries. One important class are permutation symmetries, i.e., the neurons or channels in the same layer can be randomly permuted and, if the incoming and outgoing connections are adjusted accordingly, form structurally different, yet functionally identical networks. An example is given in Sec. 2. A layer with $n$ neurons has $n!$ permutations. A deep network with $l$ such layers has $\prod_{l}n!$ functionally identical solutions. For example, Ainsworth et al. (2022) calculated that ResNet50 contains $10^{55^{\prime}109}$ permutation symmetries.
@@
We leverage permutation invariance of neural networks to permute the units of a layer in descending order based on their importance scores. This operation doesn’t change the subnetwork function, but ensures that the nested subnetworks keep the most important units of each layer as part of their architecture, thus simplifying each subnetwork architecture search. To preserve the correct feature extraction process, the permutation operation is performed layerwise and the structural elements to permute depend on the layer type. In a standard convolution layer, for example, the weights are stored as a four-dimensional tensor (see Fig. 2). When the convolutional filters are permuted, the kernels in the subsequent layer are similarly reordered to maintain consistent input channel and kernel sequence during the forward pass. This permutation process involves manipulating the original weight tensors, initially unfolding the layer’s tensor into a series of three-dimensional tensors–representing the convolutional filters–and subsequently unrolling each filter’s tensor into a set of two-dimensional tensors, corresponding to the filter’s kernels.
3.2 Importance and cost measures
@@
How do we identify key neurons or convolutional filters essential for constructing a high-precision subnetwork? Several studies employ performance scores to assess the significance of weights, neurons, and channels within a neural network. NestDNN (Fang et al., 2018) ranks each convolutional filter by calculating the L2 distance between feature maps of the same class and those of different classes. This method uses all the training data, which is often unavailable due to resource constraints when the model is used in production. Magnitude-based pruning methods, that use the magnitude of the weights to estimate their importance (Han et al., 2015; Zhu and Gupta, 2017; Cai et al., 2020), provide a computationally efficient way to rank the model’s computational units. However, numerous studies (Molchanov et al., 2019; Mozer and Smolensky, 1988) have reported that the magnitude does not necessarily reflect the importance of a unit in terms of its contribution to the loss reduction.
@@
In REDS, we adopt an efficient approach to score each model’s computational units. This approach can be used on hardware platforms, which support on-device training through gradient descent. We compute for each of the encoder’s weight $i$ the importance as $I_{i}=|g_{i}\gamma_{i}|$, where $g_{i}$ is the sum of the accumulated gradients computed from backpropagation (Rumelhart et al., 1986) and $\gamma_{i}$ is the scalar value of weight $i$ (Molchanov et al., 2019; Shen et al., 2022). For each computational unit $c$, such as a convolutional filter or a fully-connected neuron, let $\mathcal{W}_{c}$ be the set of its weights. Then its importance score, denoted as $I_{c}$, is computed as the grouped sum of the importance scores over all these weights:
@@
The computational unit $I$ approximates the squared difference of the loss when a unit is removed, thus it quantifies the contribution of each unit to the final prediction loss (Molchanov et al., 2019). Each computational unit is characterized by an importance score, denoted as $I_{c}$, and computational costs defined in terms of model latency and peak memory usage. For each unit we compute an inference latency and peak memory usage predictors on a device. The former is defined as the number of multiply–accumulate operations (MACs) (Liberis et al., 2021), while the latter is computed as the size in bytes of the activation map.

### v3

@@ 1 Introduction
Data processing pipelines in edge devices increasingly rely on deep learning models to identify patterns and extract insights from multimodal IoT data. Examples include predictive maintenance in industrial automation, object identification and tracking in smart camera systems (Qu et al., 2022), activity and healthcare trackers in mobile (Ravi et al., 2016), wearable and hearable applications (Maag et al., 2017; Sabry et al., 2022). In all these systems, deep models are deployed and run along with other tasks, under constraints and priorities dictated by the current context and available resources, including storage, CPU time, energy and bandwidth. A large body of work explores different techniques to optimize and compress deep models without hurting accuracy and generalization abilities, while accelerating their execution in software (Boehm et al., 2018) and in hardware (Jouppi et al., 2017). Model pruning and quantization (Han et al., 2016) have become part of standard deep learning deployment pipelines, e.g., TFLMicro (David et al., 2021) and TensorRT (Vanholder, 2016), to enable deep learning on severely constrained embedded hardware operated by low-power microcontrollers with only a few kB of RAM. Pruning covers a set of methods that reduce the model size by eliminating unimportant operations (weights, neurons, kernels) in the model (Dai et al., 2018; Li et al., 2017). These methods date back to the optimal brain damage (LeCun et al., 1989) and the optimal brain surgeon (Hassibi and Stork, 1993), which suggest to prune the weights based on the Hessians of the loss function. Recent pruning methods (Han et al., 2015; Entezari and Saukh, 2019; Timpl et al., 2021; Corti et al., 2022) propose to prune network weights with a low magnitude or a low magnitude increase. (Li et al., 2017) propose to prune channels in CNNs based on a filter weight norm, while Hu et al. (2016) use the average percentage of zeros in the output to prune unimportant channels.
@@
One drawback of compile-time optimizations, e.g., pruning, is that the resulting models are resource-agnostic. They thus yield suboptimal performance in many interesting applications, where resource availability depends on different dynamically changing factors such as available energy, task priority and timing constraints. Another drawback of one-shot model compression techniques, is that these are applied to the whole model, making exploration of different options for a resource-aware on-device model reconfiguration challenging. Both problems are described in detail below.
@@
Dynamic resource constraints. Many interesting applications can make use of resource-aware deep models, i.e., models that can adapt their execution to available computational resources and time constraints. For example, camera image processing by a drone or a car may depend on the respective speed (Qu et al., 2022). Processing high-value data can justify using more energy and computational time than when running regular environment scans (Gherman et al., 2021). A naive solution to address dynamic resource constraints is to store several independent deep models and switch between them as resource availability and task priorities change. The drawback of this approach is the increased memory consumption to store these independent models, which does not scale, and the overhead of switching between models at runtime, e.g., by loading these from flash to RAM and reallocating the necessary tensors.
@@
Several approaches have been proposed to adapt a model to dynamic resource constraints. These methods can be classified into two groups: those that implement early exit predictions at a cost of a reduced accuracy, e.g., BudgetRNN (Kannan and Hoffmann, 2021) and ASVN (Bambusi et al., 2021), and those that build a subnetwork structure using weight sharing, with each subnetwork being more efficient yet possibly less accurate, e.g., Slimmable Subnets (Yu et al., 2019), DRESS (Qu et al., 2022) and NestDNN (Fang et al., 2018). This work falls into the latter category but leverages structured sparsity constructively to enable additional flexibility and hardware support on IoT devices.
@@
Deep learning support on IoT devices. Modern IoT hardware often provides support for running deep learning models. This support includes floating point instructions, dedicated instructions for frequent operations, such as MLA, FMA, and their vector versions. Deep neural network frameworks use hardware-specific features on supported platform to provide maximum speedup. However, the methods that introduce subnetwork structures to a model can not rely on the toolchain support to optimize the execution of each subnetwork. For example, fine-grained weight pruning may lead to accelerated execution due to an abundance of zeros in the weight matrices and sparse filters, yet unconstrained locations of these zeros present a difficulty in using unstructured sparsity. Additionally, specific algorithms are designed to effectively manage various sparsity levels, which may or may not be supported by the available hardware (Hoefler et al., 2021). To overcome the issues, NestDNN (Fang et al., 2018) prunes convolutional filters; these have to be paged in or out when switching from one multi-capacity model to another. DRESS (Qu et al., 2022) relies on specialized hardware to ensure the applied sparsity pattern translates into computation efficiency. Hardware accelerators may appear too power-hungry and expensive for battery-powered IoT devices. Specialized hardware is also inflexible as the deep learning technology advances fast and hardware modernizations are costly.
Contributions. We present the design of Resource-Efficient Deep Subnetworks (REDS), featuring a nested submodel structure to address dynamic resource constraints on mobile and IoT devices. REDS introduces a novel hardware-aware method for designing nested subnetwork structures, i.e., structured sparsity patterns, by formulating and solving an iterative knapsack problem.
This method provides both theoretical guarantees and empirical evidence of outstanding solution performance. Furthermore, we leverage permutation invariance of neurons (Entezari et al., 2022) to keep the subnetwork weight tensors in contiguous memory regions, i.e., dense layers remain dense in all lower-capacity subnetworks. Compared to standard pruning methods, which encode structural sparsity patterns using binary masks and apply uniform pruning ratios across all layers, REDS enforces that all retained neurons or filters form a single contiguous block in each weight tensor. Each REDS subnetwork can be specified by simple (offset, length) parameters without storing any masks. Moreover, REDS computes layer-specific pruning ratios and these are determined based on each layer’s contribution to the model’s representational capacity and the cost associated with application-specific computational constraints. REDS enables precise control over inference costs while preserving representational performance. In contrast to Shen et al. (2022), REDS knapsack formulation models the dependency between layers by iteratively chaining constraints to enforce each subnetwork’s weight tensor dimensions to be functionally correct (see Fig. 2). Furthermore, REDS knapsack formulation can be extended to constrain each subnetwork’s peak memory usage, i.e., activation memory usage, defined as the major memory bottleneck for enabling neural networks on the edge computing devices  (Lin et al., 2023). Our theoretical findings in the Appendix suggest that the knapsack method applied to the smallest subnetwork first and letting the REDS structure iteratively grow, i.e., the bottom-up approach, is more effective than starting from the largest subnetwork and pruning it down to the smallest one, i.e., the top-down approach. We empirically show this in Sec. 4.
REDS code is publicly available. [^1] Link: https://github.com/FraCorti/REDS_TOMC Our contributions are:
@@
  • We formulate the optimization problem as an iterative knapsack problem (Della Croce et al., 2019), define a novel generalized knapsack framework for depth-wise convolutions and vision transformers, present theoretical analysis and provide empirical evidence of the solution effectiveness, especially in the low-data regime (Sec. 3 and Sec. 4).
  • REDS uses the permutation invariance of neurons to facilitate hardware-specific optimizations (Sec. 3). In particular, for resource-constrained devices that use data caches, REDS provides a compile-time optimization to ensure all subnetwork weights reside in contiguous memory (Sec. 5).
In the next section we explain REDS on a fully-connected neural network (DNN).
2 REDS Applied to a Simple Network
@@
We first illustrate advantages of REDS on a fully-connected network depicted in Fig. 1. The three dense weight matrices $W_{1}^{3\times 4}$, $W_{2}^{4\times 4}$ and $W_{3}^{4\times 2}$ connect neurons in consecutive layers that are stored in memory as unfolded one-dimensional arrays. We assume that the row-major format is used to store and access model weights in memory, which is the standard choice on most hardware architectures. Given a trained and optimized model, weight matrices are mapped to continuous memory regions, allowing for a straightforward use of vector instructions to speed up on-device inference.
@@
REDS organizes a model into a set of nested submodels, i.e., the active weights of a child subnetwork are fully contained in its parent subnetwork. To enable this structure, we slice each parent subnetwork into an active and an inactive part, where the active part shapes the child subnetwork. REDS uses structured sparsity, i.e., model slicing occurs at the level of individual neurons and convolutional filters. Slicing off two neurons in the first hidden layer in Fig. 1 leads to removing two columns and two rows in the weight matrices $W_{1}$ and $W_{2}$, respectively.
@@
First, although any neuron can be removed from a layer, we re-order neurons to have a group of active neurons followed by inactive neurons due to the permutation invariance phenomenon of neural networks (Entezari et al., 2022). A permutation does not change the function of the network, but allows optimizing the memory layout to keep subnetwork weights in contiguous memory. The above observations make the subnetwork layout to be stored efficiently on-device. In fact, it is sufficient to store one integer value denoting the slicing point for each layer in each subnetwork, corresponding to the number of active computational units. This is possible since active and inactive units build contiguous groups in memory. Activating a particular subnetwork means changing the size of the dimension of the weight tensor in each layer. An additional bit indicates optimization for caches, i.e., that a flipped operation order is applied.
@@
Second, REDS achieves minimal adaptation overhead: only the width of the layers has to be updated to switch to a different model. Switching from one REDS submodel to another requires adjusting only the width sizes of the layers. Due to the local scope of a slice, the modification affects only the incoming and outgoing connections; inactive computational units do not participate in inference. Examples of slicing dense and convolutional networks are shown in Fig. 1 and Fig. 2. The computational cost of running inference using a subnetwork is not affected by the presence of other REDS networks, in sharp contrast to the approaches that use binary masks to select active neurons or channels Qu et al. (2022); Mishra et al. (2021).
@@
In the end, pruning a neuron in one layer may remove more multiply–accumulate operations (MACs) than in another layer. In our example in Fig. 1, removing one neuron in the first hidden layer yields more MAC reduction than removing a neuron in the second hidden layer due to a different number of connections. Moreover, the contribution of these neurons to model accuracy may be different. Related research suggests several importance measures for individual weights, neurons, and convolutional filters (Shen et al., 2022; Molchanov et al., 2019; Hoefler et al., 2021; Frantar and Alistarh, 2022). These can be used to optimize a network for performance while keeping essential elements. Our problem is different: we build a nested structure, which requires iterating over each subnetwork. Previous approaches Fang et al. (2018); Qu et al. (2022) start from the largest network or implement a top-down pruning and a bottom-up freeze-and-regrow algorithm. We formulate the optimization problem for a parent-child subnetwork pair as a variant of the knapsack problem (Shen et al., 2022). To extend the solution to multiple nested subnetworks, we generalize the approach to an iterative knapsack problem (Della Croce et al., 2019).
@@
In the following sections, we focus on specific contributions and design choices for REDS on the way to a fully-functioning framework. Sec. 3 discusses a pipeline to convert a model to REDS and presents a hardware-aware iterative knapsack method to choose the model slicing points. Sec. 5 revisits the memory layout optimization for data cache and provides an empirical evaluation of the achieved gain on embedded and mobile hardware platforms.
@@ 3 Resource-Efficient Deep Subnets
Before detailing the REDS construction pipeline, we first argue that layer slicing by skipping a contiguous sequence of computational units, whether neurons or convolutional filters, from any layer except the final classifier, preserves the representational capacity of the resulting REDS subnetworks. The pruned computational units can be reordered to form a contiguous sequence without changing the network function. We then describe REDS construction pipeline. At the core is the novel method to decide on the choice of the subnetwork structure of REDS based on the knapsack problem. Finally, we discuss how to fine-tune the submodels in REDS to recover lost accuracy due to the applied structural pruning of computational units.
@@ 3.1 Permutation invariance
The optimization landscape of neural networks contains an abundance of minima and saddle points as a result of numerous symmetries. One important class of symmetries is formed by permutations, i.e., the neurons or filters in the same layer can be randomly permuted and, if the incoming and outgoing connections are adjusted accordingly, form structurally different, yet functionally identical networks. An example is given in Sec. 2. A layer with $n$ neurons has $n!$ permutations. A deep network with $l$ such layers has $\prod_{l}n!$ functionally identical solutions (Entezari et al., 2022).
@@
We leverage permutation invariance of neural networks to permute the units of a layer in descending order based on their importance scores. This operation does not change the subnetwork function, but ensures that the nested subnetworks keep the most important units of each layer as part of their architecture. To preserve the correct feature extraction process, the permutation operation is performed layerwise and the structural elements to permute depend on the layer type. In a convolutional layer, for example, the weights are stored as a four-dimensional tensor (see Fig. 2). When the convolutional filters are permuted, the kernels in the subsequent layer are similarly reordered to maintain consistent input channel and kernel sequence during the forward pass. This permutation process involves manipulating the original weight tensors, initially reshaping the layer’s tensor into a set of three-dimensional tensors representing the convolutional filters. Each filter tensor is then reshaped into a set of two-dimensional tensors, each corresponding to one of the filter’s kernels.
3.2 Cost measures
@@
How do we identify key neurons or convolutional filters essential for constructing a high-precision subnetwork? Several studies use importance scores to quantify the significance of network weights, neurons, and channels. NestDNN (Fang et al., 2018) ranks each convolutional filter by calculating the L2 distance between feature maps of the same class and those of different classes. This method uses all the training data, which is often unavailable due to resource constraints when the model is deployed on edge devices. Magnitude-based pruning methods, that use the magnitude of the weights to estimate their importance (Han et al., 2015; Zhu and Gupta, 2017; Cai et al., 2020), provide a computationally efficient way to rank the model’s computational units. However, several studies (Molchanov et al., 2019; Mozer and Smolensky, 1988) have reported that the magnitude does not necessarily reflect the importance of a unit in terms of its contribution to the loss reduction.
@@
In REDS, we compute for each of the encoder’s weight $i$ the importance as $I_{i}=|g_{i}\gamma_{i}|$, where $g_{i}$ is the sum of the accumulated gradients computed from backpropagation (Rumelhart et al., 1986) and $\gamma_{i}$ is the scalar value of weight $i$ (Shen et al., 2022). Weights with larger accumulated gradients $g_{i}$ contribute more to reducing the cross-entropy loss function, making them critical for model’s accuracy preservation (Molchanov et al., 2019). For each computational unit $c$, such as a convolutional filter or a fully-connected neuron, let $\mathcal{W}_{c}$ be the set of its weights. Then its importance score, denoted as $I_{c}$, is computed as the grouped sum of the importance scores over all these weights:
@@
The computational unit $I$ approximates the squared difference of the loss when a unit is removed; thus it quantifies the contribution of each unit to the model’s accuracy (Molchanov et al., 2019). Each computational unit is characterized by an importance score, denoted as $I_{c}$, and computational costs defined in terms of model latency and memory usage. We compute inference latency and memory usage predictors for each unit on an edge device. Inference latency is defined as the number of multiply–accumulation operations (MACs) (Liberis et al., 2021), while memory usage is calculated as the size in bytes of the output tensor.
REDS adopts the gradient-based importance score formulation introduced by (Molchanov et al., 2019) reported in Eq. (1). Related unit scoring methods such as SNIP (Lee et al., 2018), which computes filter importance scores using a similar gradient-based metric, and OSTR (Yang et al., 2024), which derives operation-strength scores from the architecture-parameter gradients during neural architecture search, have also been proposed. However, a detailed comparison with these methods is beyond the scope of this work and is left for future investigation.
3.3 Bottom-up (BU) and top-down (TD) heuristics
REDS examines two heuristics for solving the iterative knapsack problem named bottom-up (BU) and top-down (TD). The former iteratively calculates the subnetwork architectures by considering the tightest constraints for the smallest subnetwork first. Once a solution is found by the knapsack solver, these units are frozen, i.e., they are now part of all nested subnetworks, and the second smallest subnetwork is being computed by the solver. The latter top-down method determines solutions by considering the weakest constraints first and then iteratively searching the architectures for increasingly smaller subnetworks. Related work Cai et al. (2020) uses a variant of the top-down approach, i.e., progressive shrinking. Our theoretical analysis of the iterative knapsack problem (based on the classical 0-1 knapsack problem) in the Appendix shows that the bottom-up approach promises a better worst-case performance. In particular, we prove that the solution found by the two-stage bottom-up iterative knapsack heuristic is not worse than $\frac{2}{3}\cdot Opt$, where $Opt$ is the optimal solution of the knapsack with full capacity, and that this bound is tight. We then show that the tight bound for the top-down iterative knapsack is $\frac{1}{2}\cdot Opt$, where $Opt$ in this case is the optimal solution of the knapsack with reduced capacity. Since our generalized problem suited for depthwise separable convolutional architectures has the classical 0-1 knapsack as its core problem, we believe that a similar result is valid for this case as well.